OpenStudy (anonymous):

Help! just see if you could?! The number of State Parks in California is 36 more than one third the number of State Parks in Florida. There are 85 State Parks in California. Which of the following equations would you use to find the number of State Parks in Florida? Note: f represents the number of state parks in Florida. 3(f − 36) = 85 1 over 3f + 36 = 85 Quantity f minus 36 all over 3 equals 85 Quantity f plus 36 all over 3 equals 85

OpenStudy (countryboy15):

what do you think???

OpenStudy (anonymous):

"36 more than" <<<--- + 36 1/3 of Florida <<<--- (1/3)F 85 = (1/3)F + 36 Solve for F: 3 ( 85-36) = F

OpenStudy (countryboy15):

"36 more than" <<<--- + 36 1/3 of Florida <<<--- (1/3)F 85 = (1/3)F + 36 Solve for F: 3 ( 85-36) = F

OpenStudy (anonymous):

Umm what does that all mean exactly??

OpenStudy (anonymous):

The number of State Parks in California is 36 more than one third the number of State Parks in Florida. 85 is 36 more than one third the number of State Parks in Florida. 85 = 36 + one third the number of State Parks in Florida 85 = 36 + f/3 this is the equation I would use. Of course the equation can be simplified and solved, and you are not showing us your list of choices. If I were to continue this problem, I would subtract 36 from both sides and get another equation, 49 = f/3 f = 147

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

Yes that does, I think I will read it over a couple times just to make sure I got it, lol.

OpenStudy (countryboy15):

it means that the answers c need keeps beting me to the point lol

OpenStudy (anonymous):

lolz thx guys do I get medal?

OpenStudy (anonymous):

Yeah I got it now. Thank you both very much!!!

OpenStudy (anonymous):

yw! Glad to help

OpenStudy (countryboy15):

np also need please medal me

OpenStudy (countryboy15):

there passed the medal so we all got one

OpenStudy (anonymous):

click bestresponse

OpenStudy (countryboy15):

tag use if need help again thanks

OpenStudy (anonymous):

Thanks again!